OpenStudy (anonymous):

What is the LCD. 1/2 and 3/14

OpenStudy (anonymous):

take common

OpenStudy (anonymous):

what is full form of LCD

OpenStudy (anonymous):

Lowest Common Divisior

OpenStudy (anonymous):

oh thanks sheg

OpenStudy (anonymous):

no its denominator.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Yeah.

OpenStudy (lilg132):

lowest common denominator your right

OpenStudy (anonymous):

Find LCD. Build the fractions to equivalent fractions aving the LCD as the denominator.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

still confused what is LCD

OpenStudy (anonymous):

nevermind.

OpenStudy (anonymous):

no reason to worry

OpenStudy (anonymous):

can u divide 3/14 by 1/2??

OpenStudy (anonymous):

\[\frac{\frac{3}{14}}{\frac{1}{2}}\]

OpenStudy (anonymous):

oh! i got it............